# Self-similar sets and Lipschitz graphs

### v1

@@ 5.2. The generic graph construction algorithm
$E^{m}_{n}=\bigsqcup_{i=1}^{\left(2^{2m-1}\right)^{n}}Q_{i}^{n}$
where $Q_{i}^{n}\prec Q_{i+1}^{n}$ for each $i\in\left\{1,\ldots,\left(2^{2m-1}\right)^{n}-1\right\}$. Images of some $E^{m}_{n}$ are in Figure 5.
@@
To apply Proposition 4.1, we check that $\left\{E_{n}^{m}\right\}_{n=1}^{\infty}$ satisfies the hypotheses. As in the proof of Proposition 5.2, $\left\{E_{n}^{m}\right\}_{n=1}^{\infty}$ is nested, each $E_{n}^{m}\subseteq C_{mn}$ is a finite, disjoint union of cubes, each of sidelength $4^{-mn}$. The bound (18) holds with $\lambda=\sqrt{2}$, and (20) holds with $c=\sqrt{2}$ and $\sigma=4^{-m}$.
@@ 6.1. Favard length
Recall that the function ||$\cdot$|| ⋅ | from the $\sigma$-algebra of Lebesgue measurable sets in $\mathbb{R}$ to $[0,\infty)$ represents the Lebesgue measure.
@@
[…]
@@
[…]
@@
$I\subseteq\bigcup_{\ell=1}^{M}\left(3+\frac{\varepsilon}{\delta}\right)J_{\ell}.$
@@
$\bigsqcup_{\ell=1}^{M}\frac{1}{3}J_{\ell}\subseteq\bigcup_{k=m_{0}}^{N^{n}}I_{k}\subseteq P_{\theta}(C_{n})$
@@
$\frac{1}{3}\sum_{\ell=1}^{M}\left|J_{\ell}\right|=\left|\bigsqcup_{\ell=1}^{M}\frac{1}{3}J_{\ell}\right|\leq\left|P_{\theta}\left(C_{n}\right)\right|.$
@@
$\left|P_{\theta}(C_{n}(\delta_{n}))\right|\leq\left|\bigcup_{k=1}^{m_{0}-1}I_{k}(\delta_{n})\right|+\left|\bigcup_{k=m_{0}}^{N^{n}}I_{k}(\delta_{n})\right|\leq 10\left|P_{\theta}\left(C_{n}\right)\right|.$
@@ 6.2. Extracting a sub-IFS and building the graph
For each $\ell\in\left\{1,\ldots,M\right\}$, there exists $j^{(m)}\in\left\{1,\ldots,N\right\}^{m}$ so that $J_{\ell}=I_{j^{(m)}}$. Define each $\varphi_{\ell}=f_{j^{(m)}}$ and note that $\varphi_{\ell}(x)=\rho_{\ell}x+\zeta_{\ell}$, where $\rho_{\ell}=r_{j_{1}}\ldots r_{j_{m}}$. With $E:=\bigcup_{\ell=1}^{M}\varphi_{\ell}(K)\subseteq C_{m}$, we see that $P_{\theta}\left(E\right)=\bigcup_{\ell=1}^{M}P_{\theta}\left(\varphi_{\ell}(K)\right)=\bigsqcup_{\ell=1}^{M}J_{\ell}$ so that, by construction, $P_{\theta}(E)$ is a union of $M$ $\delta_{m}$-separated intervals.
@@
Now we check that $\left\{E_{n}\right\}_{n=1}^{\infty}$ satisfies the hypotheses of Proposition 4.1. By construction, $\left\{E_{n}\right\}_{n=1}^{\infty}$ is nested and each $E_{n}:=\bigsqcup_{j=1}^{M^{n}}K^{n}_{j}$ consists of $M^{n}$ connected components, each a translated and rescaled copy of $K$. It follows that (18) holds with $\lambda={\operatorname{diam}(K)}\nu^{-1}$. For each $n,j$, $\operatorname{diam}\left(K_{j}^{n}\right)\leq\operatorname{diam}(K){r_{N}}^{mn}$ and then (20) holds with $c=\operatorname{diam}(K)$, $\sigma=r_{N}^{m}<1$.
@@
$\dim\left(\Gamma\cap C\right)\geq\dim\left(E\right)\geq 1-\varepsilon.$
@@ 7.1. Reduction to a uniform iterated function system
Now we to check that $[\kappa_{1},\kappa_{2}]\cap\mathbb{N}\neq\varnothing$. By the Mean Value Theorem, there exists $\bar{\kappa}\in(\kappa_{1},\kappa_{2})$ so that
@@
Let $\left\{f_{k}\right\}_{k=1}^{M}$ be an IFS that satisfies the open set condition, has attractor $H$ with $\dim(H)=1$, Ahlfors lower and upper constants $a$ and $b$, respectively, and overlapping index $\omega$.
If for some $\kappa\in\mathbb{N}$, $\left\{\varphi_{j}\right\}_{j=1}^{N}\subseteq\left\{f_{k^{(\kappa)}}\right\}$ is a UIFS with scale factor $r$, attractor $C\subseteq H$, and $\gamma:=\dim(C)<1$, then for all $x\in C$ and all $\rho\in(0,1)$, it holds that
(36) $\mathcal{H}^{\gamma}\left(C\cap B(x,\rho)\right)\leq\frac{8\omega b\operatorname{diam}(K)}{a\nu(K)}r^{\gamma-1}\rho^{\gamma},$
@@
(37) $\mathcal{H}^{\gamma}(C)\geq\frac{a\nu(K)}{8\omega b\left[\operatorname{diam}(K)\right]^{1-\gamma}}r^{1-\gamma},$
@@
Since $\omega$ is the overlapping index of $\left\{f_{k}\right\}_{k=1}^{M}$, then by Lemma 3.14, the overlapping index of the generations $\left\{C_{n}\right\}_{n=1}^{\infty}$ is at most $\omega$. It follows that $\left\{B_{j^{(n)}}\right\}$ can be partitioned into $\omega$ subsets, each of which contains disjoint convex sets.
@@
Then we see that
$\frac{\nu(K)}{\operatorname{diam}(K)}r\rho\cdot\#\mathcal{S}(x,\rho)$ $\leq\frac{\nu(K)}{\operatorname{diam}(K)}\sum_{j^{(k)}\in\mathcal{S}(x,\rho)}\operatorname{diam}\left(K_{j^{(k)}}\right)\leq 4\sum_{j^{(k)}\in\mathcal{S}(x,\rho)}\text{rad}\left(B_{j^{(k)}}\right)$
$\leq\frac{4}{a}\sum_{j^{(k)}\in\mathcal{S}(x,\rho)}\mathcal{H}^{1}\left(H\cap B_{j^{(k)}}\right)\leq\frac{4\omega}{a}\mathcal{H}^{1}\left(H\cap\bigcup_{j^{(k)}\in\mathcal{S}(x,\rho)}B_{j^{(k)}}\right)$
$\leq\frac{4\omega}{a}\mathcal{H}^{1}\left(H\cap B(x,2\rho)\right)\leq\frac{8\omega b}{a}\rho,$
where we recall that $a$ and $b$ are the Ahlfors lower and upper constants of $H$. Therefore, #$\#\mathcal{S}(x,\rho)\leq\frac{8\omega b\operatorname{diam}(K)}{a\nu(K)}r^{-1}$, and an application of Jensen’s inequality shows that
(38) $\sum_{j^{(k)}\in\mathcal{S}(x,\rho)}\left[\operatorname{diam}\left(K_{j^{(k)}}\right)\right]^{\gamma}$$\leq{\#\mathcal{S}(x,\rho)}^{1-\gamma}\left[\sum_{j^{(k)}\in\mathcal{S}(x,\rho)}\operatorname{diam}\left(K_{j^{(k)}}\right)\right]^{\gamma}$
$\leq\frac{8\omega b\operatorname{diam}(K)}{a\nu(K)}r^{\gamma-1}\rho^{\gamma}.$
@@
$\sum_{j_{k+1}=1}^{N}\left[\operatorname{diam}(K_{(j^{(k)},j_{k+1})})\right]^{\gamma}=\sum_{j=1}^{N}\left[r\operatorname{diam}(K_{(j^{(k)})})\right]^{\gamma}=\left[\operatorname{diam}(K_{(j^{(k)})})\right]^{\gamma},$
@@
$\frac{8\omega b\operatorname{diam}(K)}{a\nu(K)r^{1-\gamma}}\sum_{i=1}^{L}\operatorname{diam}\left(A_{i}\right)^{\gamma}$ $=\frac{8\omega b\operatorname{diam}(K)}{a\nu(K)r^{1-\gamma}}\sum_{i=1}^{L}\rho_{i}^{\gamma}\geq\sum_{i=1}^{L}\sum_{j^{(k_{i})}\in\mathcal{S}(x_{i},\rho_{i})}\left[\operatorname{diam}\left(K_{j^{(k_{i})}}\right)\right]^{\gamma}$
@@ 7.2. Existence of large projections
    if $\phi\in J$, then there exists a subcollection $\mathcal{Q}_{\phi}$ of $\mathcal{Q}$ of cardinality at least $\varepsilon\delta\#\mathcal{Q}$ such that the orthogonal projections of the sets in $\mathcal{Q}_{\phi}$ onto a line with direction $\phi$ are all disjoint and $\rho$-separated;
@@
Let $\left\{\varphi_{j}\right\}_{j=1}^{N}$ be an IFS that satisfies the open set condition, has uniform scale factor $r$, similarity dimension $\gamma\in\left(0,1\right)$, attractor $C$, and upper Ahlfors regularity constant $b_{0}$. Let $K$ be a compact, convex, $\nu$-non-degenerate set for which $C\subseteq K$. Let $\left\{C_{n}\right\}_{n=0}^{\infty}$ denote the generations of $\left\{\varphi_{j}\right\}_{j=1}^{N}$ with respect to $K$, and let $\omega_{0}$ denote the overlapping index of the generations.
@@
    If $\phi\in J$, then there exists a subset $C_{n,\phi}\subseteq C_{n}$, with $N_{\phi}\geq\varepsilon\delta_{0}r^{-n\gamma}$ connected components, for which $P_{\phi}(C_{n,\phi})$ is a disjoint union of $N_{\phi}$ $r^{n}$-separated intervals;
The constant $\delta_{0}$ given in (43) depends on $\gamma$, $\nu(K)$, $\operatorname{diam}(K)$, $H^{\gamma}(C)$, $b_{0}$, $\omega_{0}$, and is independent of $n$.
@@
Fix $n\in\mathbb{N}$. We have $C_{n}=\bigcup_{j^{(n)}}K_{j^{(n)}}$, where $K_{j^{(n)}}:=\varphi_{j^{(n)}}(K)$. The claimed result follows from an application of Proposition 7.4 with $\rho=r^{n}$ and a choice of $\mathcal{Q}\subseteq\left\{K_{j^{(n)}}\right\}$ that satisfies the hypotheses. We first specify $\mathcal{Q}$. Since the generations $\{C_{n}\}_{n=0}^{\infty}$ have overlapping index $\omega_{0}$, then the index set can be partitioned as $\left\{1,\ldots,N\right\}^{n}=\bigsqcup_{k=1}^{\omega_{0}}\mathcal{T}_{k}$, where the sets in each collection $\left\{K_{j^{(n)}}:j^{(n)}\in\mathcal{T}_{k}\right\}$ have disjoint interiors. Choose a maximal $\mathcal{T}\in\left\{\mathcal{T}_{k}\right\}_{k=1}^{\omega_{0}}$ in the sense that #$\#\mathcal{T}=\max\left\{\#\mathcal{T}_{k}:k\in\left\{1,\ldots,\omega_{0}\right\}\right\}$ and define
@@
By the pigeonhole principle, #$\#\mathcal{T}\geq\omega_{0}^{-1}N^{n}$. Since the similarity dimension of $C$ is $\gamma$, then $Nr^{\gamma}=1$ so that $N^{n}=r^{-n\gamma}$. Therefore, with $\alpha_{1}=\omega_{0}$,
(41) #$\#\mathcal{Q}\geq\alpha_{1}^{-1}\left(r^{-n}\right)^{\gamma}.$
@@
Assume that $B(x,\left(1+2\alpha_{0}\right)\ell)\cap C=\varnothing$. Since $C_{n}\subseteq C\left(2\alpha_{0}r^{n}\right)\subseteq C\left(2\alpha_{0}\ell\right)$, then $B(x,\ell)\cap C_{n}=\varnothing$ which implies that #$\#\mathcal{T}(x,\ell)=0$. Therefore, there is no loss in assuming that $x$ is a point for which $B(x,\left(1+2\alpha_{0}\right)\ell)\cap C\neq\varnothing$. Because each $K_{j^{(n)}}$ is contained in a ball of radius $\alpha_{0}r^{n}<\alpha_{0}\ell$, then $\bigcup_{j^{(n)}\in\mathcal{T}(x,\ell)}K_{j^{(n)}}\subseteq B(x,\left(1+2\alpha_{0}\right)\ell)$. And since $B(x,\left(1+2\alpha_{0}\right)\ell)\cap C\neq\varnothing$, then $\bigcup_{j^{(n)}\in\mathcal{T}(x,\ell)}K_{j^{(n)}}\subseteq B(y,(2+4\alpha_{0})\ell)$ for some $y\in B(x,\left(1+2\alpha_{0}\right)\ell)\cap C$. Therefore,
@@
#$\#\mathcal{T}(x,\ell)$ $\leq\alpha_{2}\left(\frac{\ell}{r^{n}}\right)^{\gamma},$
@@
We may now apply Proposition 7.4 to $\mathcal{Q}$ with $\rho=r^{n}$ to reach the conclusion. That is, for any $\varepsilon>0$, there exists a set $J\subseteq[0,\pi)$ for which $\left|[0,\pi)\backslash J\right|\leq\varepsilon$, and if $\phi\in J$, then there exists a subset $C_{n,\phi}\subseteq C_{n}$, that has at least $\varepsilon\delta\#\mathcal{Q}\geq\varepsilon\delta\omega_{0}^{-1}r^{-n\gamma}$ components, where we have used (41). In particular, Remark 7.6 shows that our claim holds with
(43) $\delta_{0}=\delta\omega_{0}^{-1}=\frac{\mathcal{H}^{\gamma}(C)}{15c_{e}\alpha_{0}^{4}\left(\alpha_{0}+1\right)\omega_{0}^{2}b_{0}}\frac{1-e^{-(1-\gamma)}}{(2+4\alpha_{0})^{\gamma}}.$

### v2

@@ 5.2. The generic graph construction algorithm
$E^{m}_{n}=\bigcup_{i=1}^{\left(2^{2m-1}\right)^{n}}Q_{i}^{n}$
where the sets are disjoint and $Q_{i}^{n}\prec Q_{i+1}^{n}$ for each $i\in\left\{1,\ldots,\left(2^{2m-1}\right)^{n}-1\right\}$. Images of some $E^{m}_{n}$ are in Figure 5.
@@
To apply Proposition 4.1, we check that $\left\{E_{n}^{m}\right\}_{n=1}^{\infty}$ satisfies the hypotheses. As in the proof of Proposition 5.2, $\left\{E_{n}^{m}\right\}_{n=1}^{\infty}$ is nested, each $E_{n}^{m}\subseteq C_{mn}$ is a finite union of disjoint cubes, each of side length $4^{-mn}$. The bound (18) holds with $\lambda=\sqrt{2}$, and (20) holds with $\sigma_{n}=\sqrt{2}\cdot 4^{-nm}$.
@@ 6.1. Favard length
Recall that the function $|\cdot|$ from the $\sigma$-algebra of Lebesgue measurable sets in $\mathbb{R}$ to $0,\infty)$ represents the Lebesgue measure.
@@
$\int_{F}\left|P_{\theta}(E(r))\right|d\theta\geq\frac{c_{M}}{b}\left|F\right|^{2}r^{1-s}.$
@@
$\int_{F}\left|P_{\theta}(E(r))\right|d\theta\geq\frac{c_{M}}{b}\left|F\right|^{2}\left(\log r^{-1}\right)^{-1}.$
@@
$I\subseteq\bigcup_{\ell=1}^{M}\left(3+\frac{\varepsilon}{\delta}\right)J_{\ell}.\qed$
@@
$\bigcup_{\ell=1}^{M}\frac{1}{3}J_{\ell}\subseteq\bigcup_{k=m_{0}}^{N^{n}}I_{k}\subseteq P_{\theta}(C_{n})$
@@
$\frac{1}{3}\sum_{\ell=1}^{M}\left|J_{\ell}\right|=\left|\bigcup_{\ell=1}^{M}\frac{1}{3}J_{\ell}\right|\leq\left|P_{\theta}\left(C_{n}\right)\right|.$
@@
$\left|P_{\theta}(C_{n}(\delta_{n}))\right|\leq\left|\bigcup_{k=1}^{m_{0}-1}I_{k}(\delta_{n})\right|+\left|\bigcup_{k=m_{0}}^{N^{n}}I_{k}(\delta_{n})\right|\leq 10\left|P_{\theta}\left(C_{n}\right)\right|.\qed$
@@ 6.2. Extracting a sub-IFS and building the graph
For each $\ell\in\left\{1,\ldots,M\right\}$, there exists $j^{(m)}\in\left\{1,\ldots,N\right\}^{m}$ so that $J_{\ell}=I_{j^{(m)}}$. Define each $\varphi_{\ell}=f_{j^{(m)}}$ and note that $\varphi_{\ell}(x)=\rho_{\ell}x+\zeta_{\ell}$, where $\rho_{\ell}=r_{j_{1}}\ldots r_{j_{m}}$. With $E:=\bigcup_{\ell=1}^{M}\varphi_{\ell}(K)\subseteq C_{m}$, we see that $P_{\theta}\left(E\right)=\bigcup_{\ell=1}^{M}P_{\theta}\left(\varphi_{\ell}(K)\right)=\bigcup_{\ell=1}^{M}J_{\ell}$ so that, by construction, $P_{\theta}(E)$ is a union of $M$ $\delta_{m}$-separated intervals.
@@
Remark 6.6.
If there exists $\theta\in[0,\pi]$ so that $\left|P_{\theta}(C)\right|\geq\bar{c}>0$, then for each $n\in\mathbb{N}$, $\left|P_{\theta}(C_{n})\right|\geq\bar{c}$. In this case, following the arguments in the proof of Proposition 6.5 leads to a similar result where now $s\geq s_{0}:=1-\frac{\log\bar{c}_{1}}{m\log(r_{N}^{-1})}$. If $\left\{f_{j}\right\}_{j=1}^{N}$ is a rotation-free, $\nu$-non-degenerate IFS that satisfies the open set condition with similarity dimension greater than $1$, the existence of $\theta$ and $\bar{c}$ is guaranteed. In fact, there are such IFS with similarity dimension equal to $1$. For example, for the 4-corner Cantor set, we can take $\theta=\arctan(1/2)$ and $\bar{c}=\frac{3}{\sqrt{5}}$.
@@
Now we check that $\left\{E_{n}\right\}_{n=1}^{\infty}$ satisfies the hypotheses of Proposition 4.1. By construction, $\left\{E_{n}\right\}_{n=1}^{\infty}$ is nested and each $E_{n}:=\bigcup_{j=1}^{M^{n}}K^{n}_{j}$ consists of $M^{n}$ connected components, each a translated and rescaled copy of $K$. It follows that (18) holds with $\lambda={\operatorname{diam}(K)}\nu^{-1}$. For each $n,j$, $\operatorname{diam}\left(K_{j}^{n}\right)\leq\operatorname{diam}(K){r_{N}}^{mn}$ and then (20) holds with $\sigma_{n}=\operatorname{diam}(K){r_{N}}^{mn}$.
@@
(31) $\dim\left(\Gamma\cap C\right)\geq\dim\left(E\right)\geq 1-\varepsilon.\qed$
Remark 6.8.
Following Remark 6.6, if there exists $\theta\in[0,\pi]$ so that $\left|P_{\theta}(C)\right|\geq\bar{c}>0$, then with $s\geq 1-\frac{\log\bar{c}_{1}}{m\log(r_{N}^{-1})}$, we can set $\varepsilon_{0}=\frac{\ln\bar{c}_{1}}{N_{0}\ln(r_{N}^{-1})}$ and $m=\lceil\frac{\ln\bar{c}_{1}}{\varepsilon\ln(r_{N}^{-1})}\rceil\geq N_{0}$. Then we see that
$\operatorname{Lip}(\Gamma)\leq\frac{\operatorname{diam}(K)}{\nu\left(\frac{r_{N}}{4N}\right)^{m}}$ $=\frac{\operatorname{diam}(K)}{\nu}\exp\left(c_{0}\varepsilon^{-1}\right).$
In comparison to the conclusion of Theorem 6.7, this is an improvement. Note that this bound is similar to the one given in Proposition 5.3 for our generic graph construction for the 4-corner Cantor set.
@@ 7.1. Reduction to a uniform iterated function system
Now we to check that $\kappa_{1},\kappa_{2}]\cap\mathbb{N}\neq\varnothing$. By the Mean Value Theorem, there exists $\bar{\kappa}\in(\kappa_{1},\kappa_{2})$ so that
@@
Let $\left\{f_{k}\right\}_{k=1}^{M}$ be an IFS that satisfies the open set condition, has attractor $H$, Ahlfors lower and upper constants $a$ and $b$, respectively, and overlapping index $\tau$.
If for some $\kappa\in\mathbb{N}$, $\left\{\varphi_{j}\right\}_{j=1}^{N}\subseteq\left\{f_{k^{(\kappa)}}\right\}$ is a UIFS with scale factor $r$, attractor $C\subseteq H$, and $\gamma:=\dim(C)<\dim(H)$, then for all $x\in C$ and all $\rho\in(0,1)$, it holds that
(37) $\mathcal{H}^{\gamma}\left(C\cap B(x,\rho)\right)\leq\frac{\tau b}{a}\left[\frac{8\operatorname{diam}(K)}{\nu(K)}\right]^{\dim(H)}r^{\gamma\dim(H)}\rho^{\gamma},$
@@
(38) $\mathcal{H}^{\gamma}(C)\geq\frac{a}{\tau b}\left[\frac{\nu(K)}{8}\right]^{\dim(H)}\left[\frac{r}{\operatorname{diam}(K)}\right]^{\dim(H)-\gamma},$
@@
Since $\tau$ is the overlapping index of $\left\{f_{k}\right\}_{k=1}^{M}$, then by Lemma 3.14, the overlapping index of the generations $\left\{C_{n}\right\}_{n=1}^{\infty}$ is at most $\tau$. It follows that $\left\{B_{j^{(n)}}\right\}$ can be partitioned into $\tau$ subsets, each of which contains disjoint convex sets.
@@
Setting $\dim(H)=s$, we see that
$\left(r\rho\right)^{s}\cdot\#\mathcal{S}(x,\rho)$ $\leq\sum_{j^{(k)}\in\mathcal{S}(x,\rho)}\left[\operatorname{diam}\left(K_{j^{(k)}}\right)\right]^{s}\leq\left[\frac{4\operatorname{diam}(K)}{\nu(K)}\right]^{s}\sum_{j^{(k)}\in\mathcal{S}(x,\rho)}\left[\text{rad}\left(B_{j^{(k)}}\right)\right]^{s}$
$\leq\left[\frac{4\operatorname{diam}(K)}{\nu(K)}\right]^{s}\sum_{j^{(k)}\in\mathcal{S}(x,\rho)}\frac{\mathcal{H}^{s}\left(H\cap B_{j^{(k)}}\right)}{a}$
$\leq\left[\frac{4\operatorname{diam}(K)}{\nu(K)}\right]^{s}\frac{\tau}{a}\mathcal{H}^{s}\left(H\cap\bigcup_{j^{(k)}\in\mathcal{S}(x,\rho)}B_{j^{(k)}}\right)$
$\leq\left[\frac{4\operatorname{diam}(K)}{\nu(K)}\right]^{s}\frac{\tau}{a}\mathcal{H}^{s}\left(H\cap B(x,2\rho)\right)\leq\frac{\tau b}{a}\left[\frac{8\operatorname{diam}(K)}{\nu(K)}\right]^{s}\rho^{s},$
where we recall that $a$ and $b$ are the Ahlfors lower and upper constants of $H$. In particular,
$\mathcal{S}(x,\rho)\leq\frac{\tau b}{a}\left[\frac{8\operatorname{diam}(K)}{\nu(K)}\right]^{s}r^{-{s}}$
and an application of Jensen’s inequality shows that
(39) $\sum_{j^{(k)}\in\mathcal{S}(x,\rho)}\left[\operatorname{diam}\left(K_{j^{(k)}}\right)\right]^{\gamma}$ $\leq{\#\mathcal{S}(x,\rho)}^{1-\frac{\gamma}{s}}\left(\sum_{j^{(k)}\in\mathcal{S}(x,\rho)}\left[\operatorname{diam}\left(K_{j^{(k)}}\right)\right]^{s}\right)^{\frac{\gamma}{s}}$
$\leq\frac{\tau b}{a}\left[\frac{8\operatorname{diam}(K)}{\nu(K)}\right]^{s}r^{\gamma-{s}}\rho^{\gamma}.$
@@
$\sum_{j_{k+1}=1}^{N}\left[\operatorname{diam}(K_{(j^{(k)},j_{k+1})})\right]^{\gamma}=\sum_{j=1}^{N}\left[r\operatorname{diam}(K_{j^{(k)}})\right]^{\gamma}=\left[\operatorname{diam}(K_{j^{(k)}})\right]^{\gamma},$
@@
$\frac{\tau b}{a}\left[\frac{8\operatorname{diam}(K)}{\nu(K)}\right]^{s}r^{\gamma-{s}}\sum_{i=1}^{L}\operatorname{diam}\left(A_{i}\right)^{\gamma}$ $=\frac{\tau b}{a}\left[\frac{8\operatorname{diam}(K)}{\nu(K)}\right]^{s}r^{\gamma-{s}}\sum_{i=1}^{L}\rho_{i}^{\gamma}$
$\geq\sum_{i=1}^{L}\sum_{j^{(k_{i})}\in\mathcal{S}(x_{i},\rho_{i})}\left[\operatorname{diam}\left(K_{j^{(k_{i})}}\right)\right]^{\gamma}$
@@ 7.2. Existence of large projections
    if $\phi\in J$, then there exists a subcollection $\mathcal{Q}_{\phi}$ of $\mathcal{Q}$ of cardinality at least $\varepsilon\delta\#\mathcal{Q}$ such that the orthogonal projections of the sets in $\mathcal{Q}_{\phi}$ onto a line with direction $\phi$ are all $\rho$-separated;
@@
Let $\left\{\varphi_{j}\right\}_{j=1}^{N}$ be an IFS that satisfies the open set condition, has uniform scale factor $r$, similarity dimension $\gamma\in\left(0,1\right)$, attractor $C$, and upper Ahlfors regularity constant $b_{0}$. Let $K$ be a compact, convex, $\nu$-non-degenerate set for which $C\subseteq K$. Let $\left\{C_{n}\right\}_{n=0}^{\infty}$ denote the generations of $\left\{\varphi_{j}\right\}_{j=1}^{N}$ with respect to $K$, and let $\tau_{0}$ denote the overlapping index of the generations.
@@
    If $\phi\in J$, then there exists a subset $C_{n,\phi}\subseteq C_{n}$, with $N_{\phi}\geq\varepsilon\delta_{0}r^{-n\gamma}$ connected components, for which $P_{\phi}(C_{n,\phi})$ is a union of $N_{\phi}$ $r^{n}$-separated intervals;
The constant $\delta_{0}$ given in (44) depends on $\gamma$, $\nu(K)$, $\operatorname{diam}(K)$, $H^{\gamma}(C)$, $b_{0}$, $\tau_{0}$, and is independent of $n$.
@@
Fix $n\in\mathbb{N}$. We have $C_{n}=\bigcup_{j^{(n)}}K_{j^{(n)}}$, where $K_{j^{(n)}}:=\varphi_{j^{(n)}}(K)$. The claimed result follows from an application of Proposition 7.4 with $\rho=r^{n}$ and a choice of $\mathcal{Q}\subseteq\left\{K_{j^{(n)}}\right\}$ that satisfies the hypotheses. We first specify $\mathcal{Q}$. Since the generations $\{C_{n}\}_{n=0}^{\infty}$ have overlapping index $\tau_{0}$, then the index set can be partitioned as $\left\{1,\ldots,N\right\}^{n}=\bigcup_{k=1}^{\tau_{0}}\mathcal{T}_{k}$, where the sets in each collection $\left\{K_{j^{(n)}}:j^{(n)}\in\mathcal{T}_{k}\right\}$ have disjoint interiors. Choose a maximal $\mathcal{T}\in\left\{\mathcal{T}_{k}\right\}_{k=1}^{\tau_{0}}$ in the sense that $\mathcal{T}=\max\left\{\#\mathcal{T}_{k}:k\in\left\{1,\ldots,\tau_{0}\right\}\right\}$ and define
@@
By the pigeonhole principle, $\mathcal{T}\geq\tau_{0}^{-1}N^{n}$. Since the similarity dimension of $C$ is $\gamma$, then $Nr^{\gamma}=1$ so that $N^{n}=r^{-n\gamma}$. Therefore, with $\alpha_{1}=\tau_{0}$,
(42) $\mathcal{Q}\geq\alpha_{1}^{-1}\left(r^{-n}\right)^{\gamma}.$
@@
Assume that $B(x,\left(1+2\alpha_{0}\right)\ell)\cap C=\varnothing$. Since $C_{n}\subseteq C\left(2\alpha_{0}r^{n}\right)\subseteq C\left(2\alpha_{0}\ell\right)$, then $B(x,\ell)\cap C_{n}=\varnothing$ which implies that $\mathcal{T}(x,\ell)=0$. Therefore, there is no loss in assuming that $x$ is a point for which $B(x,\left(1+2\alpha_{0}\right)\ell)\cap C\neq\varnothing$. Because each $K_{j^{(n)}}$ is contained in a ball of radius $\alpha_{0}r^{n}<\alpha_{0}\ell$, then $\bigcup_{j^{(n)}\in\mathcal{T}(x,\ell)}K_{j^{(n)}}\subseteq B(x,\left(1+2\alpha_{0}\right)\ell)$. And since $B(x,\left(1+2\alpha_{0}\right)\ell)\cap C\neq\varnothing$, then $\bigcup_{j^{(n)}\in\mathcal{T}(x,\ell)}K_{j^{(n)}}\subseteq B(y,(2+4\alpha_{0})\ell)$ for some $y\in B(x,\left(1+2\alpha_{0}\right)\ell)\cap C$. Therefore,
@@
$\#\mathcal{T}(x,\ell)$ $\leq\alpha_{2}\left(\frac{\ell}{r^{n}}\right)^{\gamma},$
@@
We may now apply Proposition 7.4 to $\mathcal{Q}$ with $\rho=r^{n}$ to reach the conclusion. That is, for any $\varepsilon>0$, there exists a set $J\subseteq[0,\pi)$ for which $\left|[0,\pi)\backslash J\right|\leq\varepsilon$, and if $\phi\in J$, then there exists a subset $C_{n,\phi}\subseteq C_{n}$, that has at least $\varepsilon\delta\#\mathcal{Q}\geq\varepsilon\delta\tau_{0}^{-1}r^{-n\gamma}$ components, where we have used (42). In particular, Remark 7.6 shows that our claim holds with
(44) $\delta_{0}=\delta\tau_{0}^{-1}=\frac{\mathcal{H}^{\gamma}(C)}{15c_{e}\alpha_{0}^{4}\left(\alpha_{0}+1\right)\tau_{0}^{2}b_{0}}\frac{1-e^{-(1-\gamma)}}{(2+4\alpha_{0})^{\gamma}}.$